OpenStudy (mikey666):

What are the roots of x in -10x2 + 12x − 9 = 0?

OpenStudy (asiannerd):

Do you know how to factor?

OpenStudy (asiannerd):

Try splitting up the equation and evaluate the problem

OpenStudy (sshayer):

discriminant is negative ,so it has imaginary roots.

OpenStudy (sshayer):

use quadratic formula to find the roots.

OpenStudy (holsteremission):

Or complete the square:\[\begin{align*} -10x^2+12x-9&=-10\left(x^2-\frac{6}{5}x+\frac{9}{10}\right)\\[1ex] &=-10\left(x^2-\frac{6}{5}x+\frac{9}{25}+\frac{27}{50}\right)\\[1ex] &=-10\left(x-\frac{3}{5}\right)^2-\frac{27}{5} \end{align*}\]Set equal to \(0\) and solve for \(x\).